# Forecasting Multivariate Urban Data via Decomposition and Spatio-Temporal Graph Analysis

### v1

@@ 4.1 Graph structure learning
Since the true underlying graph is unknown, we infer it from the data. To this end, we first decompose the time-series into its fundamental components and then infer the underlying graph for each component independently. This entire process is conducted offline. The graph structure is assumed to be static, hence it is not updated over time.
4.1.1 Decomposition
The first step in our approach involves decomposing the time-series into three fundamental components: trend, residual, and one or more seasonal components. This decomposition is crucial as it isolates the inherent characteristics of the time-series, thus improving the accuracy of forecasting and analysis Sohrabbeig et al. (2023). The decomposition process begins with the application of a moving weighted average filter to the time-series. The weights for this kernel are derived from the tricube function given below:
It assigns higher weights to nearby data points and lower weights to distant ones. The extracted trend is then subtracted from the original time-series to obtain the detrended series.
Next, the period of seasonality (e.g., daily or weekly patterns) must be determined. The period of seasonality can be identified using methods such as autocorrelation analysis, Fourier Transform, data visualization, or expert prior knowledge; in this work, we relied on data visualization and prior expert knowledge. To capture seasonal patterns in the detrended series, an aggregation approach with a sliding window of appropriate length and stride is used. The seasonal pattern is identified by averaging the segments covered by the sliding window. Fig. 1 illustrates this process.
Figure 1: Extracting seasonal patterns using a sliding window of size $l$ and a stride of $m$ times the length of the seasonality period, respectively. In this example, $l{=}3$ and $m{=}1$.
In each step of moving the sliding window, $S_{i}$ is calculated using the following equation:
$S_{i}=\frac{1}{l}\sum_{j=i}^{i+l-1}B_{i},$ (2)
where $B_{i}$ is the i-th block of the detrended time-series. Each block is a vector containing $p$ data points, where $p$ represents the length of the seasonality period, ranging from index $i\times p$ to $(i+1)\times p-1$. To preserve the input length, padding is added to both ends of the time-series. This padding consists of repeated copies of the first and last blocks.
Finally, the so-obtained seasonal components are subtracted from the detrended series to obtain the residual component. Each variate of the multivariate time-series, except for date and time features, undergoes this decomposition process separately. By the end of this procedure, we have extracted a trend, a residual, and one or multiple seasonal components for each variate, ensuring that the unique characteristics of each time-series are accurately captured and ready for the subsequent analysis.
4.1.2 Learning graph structure for each component
After decomposing the time-series into its constructing components, we construct the underlying graph for each component by connecting a subset of nodes. To ensure computational efficiency, each component is downsampled to a level that preserves its general patterns.
Next, the distance between two variates is measured using the Dynamic Time Warping (DTW) Müller (2007) metric. DTW is chosen for its robustness in handling nonlinear alignments, providing reliable similarity measurements between time-series. Finally, we connect each node to its $K$ nearest neighbors according to the DTW distance, with $K$ being a hyperparameter in our approach. The described procedure is outlined in Alg. 1.
Algorithm 1 Time-series Graph Construction
0:  time-series data $\{X_{i}\}_{i=1}^{D}$, seasonality periods $P_{1},P_{2},...,P_{r}$, number of nearest neighbors K
0:  Underlying Graph G
1:  Pattern Extraction:
2:  for each time-series $X_{i}$ do
3:     Extract trend component $T_{i}$ using a moving weighted average with tricube weights and then downsample
4:     Detrend the series: $\hat{X}_{i}=X_{i}-T_{i}$
5:     Extract seasonality component $S_{i}$
6:     for each seasonality period $P$ in $\{P_{1},P_{2},...,P_{r}\}$ do
7:        Apply a moving aggregation approach with window length $l\times P$ and stride $m\times P$
8:        Compute seasonality component $S_{i}$ as the average of the $n$ periods
9:     end for
10:  end for
11:  Distance Metric:
12:  Use DTW metric to calculate the distance between any two pairs of variates and construct distance matrix $D$ as shown in Fig. 3
13:  Construct the Underlying Graph
14:  Connect each node to its K nearest neighbors
Fig. 2 depicts the inferred graph from the trend component of the carbon intensity dataset, described in Section 5.1. Here the inferred graph is directional, as one variate might be dependent on the lagged version of the other one, but the opposite is not necessarily true. Fig. 3 illustrates the Dynamic Time Warping (DTW) distances between different nodes for both the trend and seasonal components of the air pollution dataset. These heatmaps reveal distinct patterns of similarity, with lower normalized DTW values indicating stronger temporal alignment between nodes.
Figure 2: The carbon intensity value is shown for a subset of states in the US at a given point in time. Dark brown (light yellow) indicates higher (lower) carbon intensity in a jurisdiction. Data has been collected from a number of power generation companies, which may supply power to different states. Consequently, the relationship between the states and energy-providing companies is many-to-many. Some arrows extend beyond the map because the image shows only a portion of the full area, and some zones connect to others not included in the display.
Figure 3: Heatmap of DTW distances between different nodes for (a) trend and (b) seasonal components of the air pollution dataset, described in Section 5.1. Note that the distance matrices are symmetrical, with lower values indicating greater closeness. All values are min-max normalized to ensure consistent scaling.
4.2 Spatio-Temporal Forecasting
@@
Now that we have learned the respective graph structure for each constituent component in an offline fashion, as described in Section 4.1.1, we proceed with the spatio-temporal forecasting step. The overall architecture of the proposed forecasting model is illustrated in Fig. 4. The input of the model is a sample of fixed-length historical values. These historical values include date and time features, along with a number of variates that we would like to predict.
@@ 4.2.1 Sample Decomposition
Figure 5: Decomposition of an electricity consumption sample over two weeks. The first subplot displays the original data alongside the sum of the trend and seasonal components, showing their closeness.
@@
First, we need to decompose an input sample, without date and time features, into its constituent components: trend, seasonality, and residual. This process follows the procedure described in Section 4.1.1, but without the need for downsampling. Fig. 5 illustrates the decomposition of a randomly selected sample from the Electricity dataset, described in Section 5.1.
4.2.2 Spatial feature extraction
@@
Our feature extraction model handles each component, that is, trend, seasonal, and residual, independently. For each component, we extract and combine three types of features, namely date-time, spatial, and temporal features.
The date-time features are processed through a linear layer to capture some temporal patterns. Simultaneously, spatial and temporal features are extracted from the input data. These temporal and spatial features are then combined and passed on to the forecasting module that uses them to generate the final output. Fig. 6 is a schematic diagram of the feature extraction module.
@@
$\hat{y}_{t}=f_{\text{forecast}}\left(f_{\text{time}}(x_{\text{time}})\oplus f_{\text{spatio-temporal}}(x)\right)$ (3)
@@
Figure 6: One layer of spatio-temporal feature extraction with time embedding. The input to this module has n channels, $C_{0}$ to $C_{n}-1$. Each component from the decomposition module—trend, seasonal, and residual—is processed independently using this layer. The spatial feature extractor is a Graph Attention Network v2 (GATv2), with a pre-inferred, static underlying graph structure that does not change over time.
Figure 7: Modeling of multivariate time-series as a graph. Each node represents a variate, and its historical data serves as the features of that node
@@
For spatial feature extraction, we utilize a model based on GATv2 Brody et al. (2021). In this model, each variate of the time-series is treated as a node, with its historical values within the sample serving as its features, as depicted in Fig. 7. GATv2 requires a connection graph to understand the relationships between different nodes (variates). The connection graph is inferred using the method described in Section 4.1.
@@
In a GNN, each node iteratively updates its state by aggregating information from its neighbors. The general formulation of a GNN layer is as follows:
$\mathbf{h}_{i}^{\prime}=\sigma\left(\mathbf{W}\mathbf{h}_{i}+\sum_{j\in\mathcal{N}_{i}}\mathbf{W}_{j}\mathbf{h}_{j}\right)$ (4)
@@
where $\mathbf{h}_{i}$ is the feature vector of node $i$, $\mathcal{N}_{i}$ represents the set of neighbors of node $i$, $\mathbf{W}$ and $\mathbf{W}_{j}$ are learnable weight matrices, and $\sigma$ is a nonlinear activation function. This aggregation allows the model to capture the dependencies between nodes and their neighbors, which is crucial for spatial feature extraction.
In GNNs, Equation 4 is used to calculate the updated node representation as a weighted sum of the neighbors’ representation. To increase learning capacity, we can incorporate this method with an attention mechanism. The Graph Attention Network (GAT) Veličković et al. (2017) enhances the GNN framework by incorporating an attention mechanism. This mechanism allows nodes to assign different amounts of importance to the neighbors, which enables capturing of changing dynamics. Specifically, the attention score $e_{ij}$ between node $i$ and its neighbor $j$ is computed as:
$e_{ij}=\text{LeakyReLU}\left(\mathbf{a}^{T}[\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}]\right)$ (5)
@@
The final output representation of node $i$ is computed as a weighted sum of its neighbors’ features:
@@
This attention mechanism allows GAT to focus on the most relevant neighbors, improving its ability to capture complex relationships in the graph.
GATv2 Brody et al. (2021), builds on the original GAT by addressing its limitations and introducing a more expressive dynamic attention mechanism. In GATv2, the attention score computation is modified to condition both the query and key nodes, enhancing the model’s flexibility. The new attention score $e_{ij}$ is defined as:
In this architecture, we employed just one layer of GATv2. The decision of whether to use one layer of GATv2 or stack more of them has been made empirically. Fig. 9 in Appendix A justifies why one layer of the GATv2 layer was employed for the carbon intensity dataset.
4.2.3 Temporal Feature Extraction
For temporal feature extraction, we utilize Temporal Convolutional Networks (TCNs) Bai et al. (2018). TCNs are designed to capture temporal dependencies in sequential data by leveraging convolution across time steps. The key components of TCNs are causal convolutions, dilations, and residual connections.
Causal convolutions ensure that the model only uses past information to predict future values, preserving the temporal order of the data. The causal convolution for a given layer $l$ is defined as:
$\mathbf{y}_{t}^{l}=\sum_{i=0}^{k-1}\mathbf{W}_{i}^{l}\mathbf{x}_{t-i}$ (9)
where $\mathbf{y}_{t}^{l}$ is the output at time $t$ for layer $l$, $\mathbf{W}_{i}^{l}$ are the convolutional filters, $\mathbf{x}_{t-i}$ are the inputs from the previous layer, and $k$ is the filter size.
Dilated convolutions expand the receptive field of the network without increasing the number of parameters, allowing the model to capture long-range dependencies. The dilation factor $d$ determines the spacing between the kernel elements. A dilated convolution is defined as:
$\mathbf{y}_{t}^{l}=\sum_{i=0}^{k-1}\mathbf{W}_{i}^{l}\mathbf{x}_{t-i\cdot d}$ (10)
where $d$ is the dilation factor, $k$ is the filter size, and the other symbols retain their previous meanings.
Residual connections help mitigate the vanishing gradient problem and facilitate the training of deeper networks. The output of each TCN layer is combined with its input through a residual connection:
$\mathbf{y}_{t}=\sigma\left(\mathbf{y}_{t}^{l}+\mathbf{x}_{t}\right)$ (11)
@@ 4.2.4 Forecasting
Once features are extracted for each component of the time-series, we employ a linear forecasting model to produce forecasts for the entire forecast horizon at once, utilizing a direct multistep forecasting approach. This choice is motivated by the demonstrated effectiveness of similar methods in previous studies Sohrabbeig et al. (2023); Zeng et al. (2023). Forecasts are generated simultaneously and independently for each component—trend, seasonal, and residual—mirroring the approach used in the feature extraction phase. Within each forecasting component, each variate is forecasted in a univariate fashion. This means that only the spatial and temporal features extracted for each specific variate in the previous step are used for forecasting that variate. Nevertheless, these features are influenced by the features of some other variates due to the way that they are generated.
The input to the forecasting module is the combination of spatio-temporal features and time embeddings. In this work, these features and embeddings are added together before being fed into the forecasting task. The decision to add rather than concatenate the time embeddings with the other extracted features was made based on our experimental results. Fig. 6 shows the input for the forecasting module.
The linear forecasting model operates on input data with the following characteristics:
  • Input Sequence Length ($L$): The length of the input time-series sequence.
  • Prediction Length ($H$): The number of future time steps to be predicted.
  • Number of Channels ($D$): The number of separate time-series channels in the input data.
The input to the model is a tensor $\mathbf{X}$ of shape $(B,D,L)$, where $B$ is the batch size, $D$ is the number of channels, and $L$ is the length of the input sequence. During the forward pass, the model computes the prediction $\mathbf{Y}$ using the input tensor $\mathbf{X}$ and the learnable parameters. This is done through a linear transformation followed by the addition of the bias term:
$\mathbf{Y}=\mathbf{W}\cdot\mathbf{X}+\mathbf{b}$ (12)
@@
5 Result
In this section, we first introduce the datasets and baseline models used for evaluation, then evaluate and compare the performance of our proposed model with the baselines in three different settings.
@@ 5.1 Datasets
For comprehensive coverage of critical factors influencing time series data, we select four urban datasets including electricity usage, weather conditions, carbon intensity, and air pollution. These datasets provide a robust benchmark for understanding the pros and cons of different time series forecasting models. More details on these datasets are provided below:
Table 1 summarizes all the main characteristics of the above datasets. With the exception of the air pollution dataset, the other datasets did not require any imputation of missing values. For the air pollution dataset, missing values were filled using the average of records corresponding to the same month, day of the week, and hour. As a preprocessing step, all columns in all datasets were standardized.
Dataset No. Features Length Time Resolution
@@
5.2 Baselines
To evaluate our model, we use three highly effective Transformer models, namely Informer Zhou et al. (2021), Autoformer Wu et al. (2021), and FEDformer Zhou et al. (2022), as our baselines. Additionally, we consider two recent state-of-the-art models, NLinear and DLinear Zeng et al. (2023). We also employed a straightforward baseline: Repeat Last, which replicates the last observation. Our next baseline is TimesFM Das et al. (2023), a foundation model for time series forecasting that leverages a decoder-only attention mechanism with input patching. TimesFM is pre-trained on a large corpus of both real-world and synthetic time series data, enabling it to achieve near state-of-the-art zero-shot performance across various datasets without the need for fine-tuning. In these experiments, TimesFM was used solely for inference and was not fine-tuned.
In this study, we did not compare our results with classical models such as ARIMA and SARIMA, due to the extensive time required for their parameter selection, which significantly slows down the training process. Instead, we based our conclusions on experimental results reported in other studies Wu et al. (2021); Zhou et al. (2021), which have shown that these classical models perform poorly on the long sequence time series forecasting (LSTF) task. Therefore, we opted not to include them in our direct comparisons.
5.3 Experiments

### v2

@@ 3.2 Graph Structure Learning
Since the underlying graph structure for each time-series component is not known in advance, we have to infer it from historical data. To this end, we first apply the decomposition process described in the previous section to the entire training dataset, then independently construct a graph for each resulting component. We note that graph structure learning is an offline task as the graph structures are assumed to be static.
The overall procedure is summarized in Algorithm 1. It starts with decomposing the time-series into its fundamental components, where the number of these components is assumed to be identical across all variates. To ensure computational efficiency, each component is first downsampled to a level that preserves its essential structural patterns. To quantify similarity between sensor readings within each time-series component, we compute pairwise distances between the variates using Dynamic Time Warping (DTW) Müller (2007). DTW is selected for its robustness in capturing nonlinear temporal alignments between time-series. Each node is then connected to its $K$ nearest neighbors according to the DTW distances, where $K$ is a hyperparameter.
Algorithm 1 Graph Structure Learning for Time-Series Components
0: time-series data $\{X_{i}\}_{i=1}^{D}$, number of nearest neighbors $K$
0: graph $G_{i}$ for each time-series component $i$
1:for each time-series variate $X_{i}$ do
2:  $\hat{X}_{i}\leftarrow\text{Detrend}(X_{i})$
3:  $\{P_{1},P_{2},...,P_{n}\}\leftarrow\text{GetSeasnonalityPeriods}(\hat{X}_{i})$
4:  for each seasonality period $P_{j}$ in $\{P_{1},P_{2},...,P_{n}\}$ do
5:   Compute seasonal component $S_{j}$
6:  end for
7:  Compute the residual component
8:end for
9:for each extracted component of $X$ denoted as $T_{i}$ do
10:  Downsample $T_{i}$
11:  Build matrix $A_{i}$ by computing DTW between variates
12:  Construct $G_{i}$ by selecting top $K$ values in $A_{i}$
13:end for
Figures 2(a) and 2(b) illustrate DTW distances between different nodes for both the trend and seasonal components of the air pollution dataset, described in Section 4.1. These heatmaps reveal distinct patterns of similarity, with lower normalized DTW values indicating stronger dependencies between nodes. Thus, the graphs constructed for the trend and seasonal components have different structures. Our ablation study, described in Section 6, confirms that this results in a significant improvement over a GNN-based forecasting model that does not perform decomposition or uses the same graph structure for all components.
Figure 2(c) depicts the graph constructed for the trend component of the carbon intensity dataset, described in Section 4.1. As it can be seen, the inferred graph contains directed edges, as one variate might be dependent on the lagged version of the other one, but the opposite is not necessarily true.
Figure 2: DTW distance heatmaps of (a) trend and (b) seasonal components of the air pollution dataset (Section 4.1); lower values indicate greater similarity, with all values min-max normalized. (c) Carbon intensity for selected US states at a given time, where darker colors indicate higher intensity. Data come from multiple power generation companies with many-to-many links to states; some arrows extend beyond the displayed map to conneced areas not shown.
3.3 Spatio-Temporal Forecasting via DST
@@
After constructing a graph for each time-series component in an offline fashion, we proceed with spatio-temporal forecasting. The architecture of the proposed forecasting model, DST, is shown in Figure 3. The model’s input is a sample of fixed-length historical values, $X^{t,l}$, and the corresponding date-time values, denoted as $X_{\text{time}}$.
@@ 3.3.1 Sample Decomposition
Figure 4: Decomposition of an electricity consumption sample over two weeks. The first subplot displays the original data alongside the sum of trend and seasonal components.
@@
We begin by decomposing an input sample, without the date-time series, into its constituent components: trend, seasonality, and residual. For decomposition, we follow the same procedure described in Section 3.1, but without downsampling. A separate forecasting model is trained for each component using the graph structure learned for that component specifically. Figure 4 illustrates the decomposition of a randomly selected sample from the Electricity dataset, described in Section 4.1. In this case, there is just one seasonal component.
3.3.2 Feature extraction
@@
The first module of DST is responsible for feature extraction. This module handles each component, i.e., trend, seasonal, and residual, separately. Specifically, for each component, we extract and then combine three types of features, namely spatial and temporal features and date-time embedding. The result is passed as input to the forecasting module, which generates the final output for that component. Figure 5 shows DST’s feature extraction and forecasting modules.
To make the date-time embedding, we form a series ($X_{\text{time}}$) by putting the timestamps of input data points in a sequence. We then pass this to an encoder that uses 1D convolution to produce the date-time embedding. Simultaneously, spatial and temporal features are extracted from the input data $X^{t,l}$. These temporal and spatial features are then combined and fed to the forecasting module as written below:
@@
$Y^{t,h}=f_{\text{forecast}}\left(f_{\text{time}}(X_{\text{time}})\parallel f_{\text{spatio-temporal}}(X^{t,l})\right),$ (3)
where $\parallel$ is the concatenation operator, $f_{\text{forecast}}$ is the forecasting model, $f_{\text{time}}$ is the date-time encoder, and $f_{\text{spatio-temporal}}$ is the spatiotemporal feature extractor, which is a combination of GATv2 and TCN as described next.
@@
Figure 5: One layer of spatio-temporal feature extraction with time embedding. The input to this module is readings from $n$ sensors, $X_{0}$ to $X_{n-1}$. Each component from the decomposition module is processed separately using this layer. The spatial feature extractor is a GATv2 that relies on the learned graph structure.
@@
For spatial feature extraction, we utilize a GNN. In this model, each variate of the time-series is treated as a node, with its historical values within the look-back window serving as its features. The GNN requires a connection graph to understand the relationships between different nodes (variates) via message passing. The connection graph is inferred using the method described in Section 3.2.
@@
In a GNN, each node iteratively updates its state by aggregating information from its neighbors. The general formulation of a GNN layer is given below
$\mathbf{h}_{i}^{\prime}=\sigma\left(\mathbf{W}\mathbf{h}_{i}+\sum_{j\in\mathcal{N}_{i}}\mathbf{W}_{j}\mathbf{h}_{j}\right),$ (4)
@@
where $\mathbf{h}_{i}$ is the feature vector of node $i$, $\mathcal{N}_{i}$ represents the set of neighbors of node $i$ in the underlying graph, $\mathbf{W}$ and $\mathbf{W}_{j}$ are learnable weight matrices, and $\sigma$ is a nonlinear activation function. This aggregation allows the model to capture the dependencies between each node and its neighbors, which is crucial for spatial feature extraction. The output $\mathbf{h}_{i}^{\prime}$ is the updated representation of node $i$, which is a weighted sum of its neighbors’ representations. To increase learning capacity, we use the attention mechanism for aggregation. Specifically, we use GATv2 Brody et al. (2021), a dynamic graph attention variant of GAT, to allow nodes to assign different amounts of importance to their neighbors in the aggregation step. This enables capturing time-varying dependecies. The attention score $e_{ij}$ between node $i$ and its neighbor $j$ is computed as:
@@
The final output representation of node $i$ is computed by passing a weighted sum of its neighbors’ features to the activation function:
@@
This attention mechanism allows the GNN to focus on the most relevant neighbors, improving its ability to capture complex relationships in the graph. We use just one GATv2 layer for spatial feature extraction.
For temporal feature extraction, we utilize Temporal Convolutional Networks (TCNs) Bai et al. (2018), which have demonstrated strong performance on sequential data across multiple domains. Our choice of TCN is motivated both by empirical results and by its architectural advantages over alternatives like LSTM and GRU.
TCNs capture temporal dependencies using convolutional operations over time steps, which offer parallelism during training, stable gradients, and a flexible receptive field. Compared to RNN-based models, such as LSTM and GRU, which process data sequentially and may struggle with long-range dependencies, TCNs enable faster training and are better suited to modeling long sequences, a key challenge in long-term time-series forecasting.
The adopted TCN has three core components Bai et al. (2018): causal convolutions, dilated convolutions, and residual connections. Causal convolutions ensure that predictions at time $t$ depend only on past inputs, preserving the temporal order.
Dilated convolutions expand the receptive field without increasing model size, enabling the network to capture long-range patterns.
Residual connections promote efficient training of deeper models and help mitigate vanishing gradients.
Overall, TCNs offer a strong balance between expressiveness, training efficiency, and robustness for long-range temporal modeling, making them especially well-suited to the demands of multivariate, long-term forecasting of urban data.
@@ 3.3.3 Forecasting
Once features are extracted for each component of the time-series, we employ a linear forecasting model to produce forecasts for the entire forecast horizon at once; this approach is commonly referred to as direct multistep forecasting. This choice is motivated by the demonstrated effectiveness of similar methods in previous studies Sohrabbeig et al. (2023); Zeng et al. (2023). Forecasts are generated simultaneously and independently for each component—trend, seasonal, and residual—mirroring the approach used in the feature extraction phase. Within each forecasting component, each variate is predicted in a univariate fashion. This means that only the spatial and temporal features extracted for each specific variate in the previous step are used for forecasting that variate. Nevertheless, these features are influenced by the features of other variates due to the way that they are extracted using a GNN.
The input to the forecasting model is the combination of spatio-temporal features and the time embedding (see Equation 3).
Figure 5 shows how the input for the forecasting model is constructed.
We adopt a linear forecasting model that predicts values in the next $h$ time steps.
The input to the model is a tensor $\mathbf{X}$ of shape $(B,D,l)$, where $B$ is the batch size, $D$ is the number of channels, and $l$ is the length of the input sequence. During the forward pass, the model computes the prediction $\mathbf{Y}$ using the input tensor $\mathbf{X}$ and the learnable parameters. This is done through a linear transformation followed by the addition of a bias term:
$\mathbf{Y}=\mathbf{W}\cdot\mathbf{X}+\mathbf{b},$ (8)
@@
3.3.4 Training
As explained earlier, DST takes advantage of a dedicated module for each component that was extracted via the decomposition method. While each module processes the respective component independently, all modules are jointly trained. For training, we used the PyTorch Lightning framework on a high-performance computing system featuring an AMD Ryzen 7 5800X 8-core processor (16 threads), 32GB RAM, and an NVIDIA GeForce RTX 3090 GPU with 24GB VRAM.
The model is optimized using the Adam optimizer with a learning rate of 0.0001 and a batch size of 32. Training is performed for a maximum of 20 epochs, with early stopping applied based on validation loss, using a patience of 3 epochs to mitigate overfitting. Gradient clipping with a threshold of 1.0 is employed to ensure training stability. The Mean Squared Error (MSE) loss function is used for training DST, while both MSE and Mean Absolute Error (MAE) are reported during evaluation to provide a comprehensive assessment of forecasting performance.
We note that the hyperparameter K, which defines the number of nearest neighbours for each node in the graph, is selected using grid search over a range spanning from 10% to 100% of all other nodes, based on validation performance.
4 Evaluation
We introduce the urban datasets and baseline models used for evaluating the long-term forecasting performance of DST.
@@ 4.1 Datasets
To comprehensively evaluate our model in realistic and diverse long-term forecasting scenarios, we utilize four multivariate time-series datasets spanning key domains: electricity usage, weather conditions, carbon intensity, and air pollution. These datasets were carefully selected to reflect a wide range of forecasting challenges, such as high dimensionality, varying temporal resolutions, and complex inter-variable dependencies. As a result, they are suitable benchmarks for assessing the generalizability and robustness of multivariate time-series forecasting models.
We describe each dataset below:
  • Electricity [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014: This dataset consists of hourly electricity consumption data from 321 customers over a period of three years. It serves as a benchmark for modeling periodic, high-dimensional energy demand, where capturing temporal patterns and seasonal cycles is critical.
  • Weather [^2] https://www.bgc-jena.mpg.de/wetter/: This dataset includes 21 meteorological metrics collected at 10-minute intervals throughout 2020 from a weather station at the Max Planck Biogeochemistry Institute. Its high-frequency nature makes it ideal for evaluating models’ sensitivity to rapid fluctuations and short-term dependencies.
  • Carbon Intensity [^3] https://www.electricitymaps.com/data-portal/united-states-of-america: This dataset contains hourly carbon intensity values recorded across 50 U.S. states and four territories over three years, this dataset introduces spatio-temporal diversity and is crucial for applications involving environmental sustainability and energy emissions forecasting.
  • Air Pollution [^4] https://archive.ics.uci.edu/dataset/381/beijing+pm2+5+data: This dataset provides hourly PM2.5 concentrations recorded by the U.S. Embassy in Beijing, along with concurrent meteorological data from the Beijing Capital International Airport Liang et al. (2015). It presents a complex, multivariate forecasting task influenced by both environmental and anthropogenic factors.
These datasets collectively span different temporal granularities, input complexities, and domain characteristics, making them an effective testbed for evaluating models intended for urban forecasting. Moreover, they align with benchmarks used in recent foundational and graph-based forecasting research Zeng et al. (2023); Wang et al. (2025), ensuring relevance and comparability. Table 1 summarizes the key characteristics of each dataset. There are no missing values in these datasets, except for the air pollution dataset, which required imputation using averages of similar temporal contexts (same month, weekday, and hour). As a preprocessing step, we standardized all features.
Dataset # Variates Length Time Resolution
@@
4.2 Baseline Forecasting Models
Following the literature review provided in Section 2, we select representative state-of-the-art baselines across multiple forecasting paradigms for comprehensive evaluation of our approach. Transformer-based models include Autoformer Wu et al. (2021) and PatchTST Nie et al. (2022), both of which are designed to effectively capture temporal patterns through attention mechanisms. Autoformer is chosen over early adaptations of transformer-based models for time-series forecasting, namely Informer Zhou et al. (2021) and Fedformer Zhou et al. (2022), primarily due to its consistently stronger performance on our datasets, making it a more suitable representative of that class of models.
From the class of linear models, we incorporate DLinear Zeng et al. (2023), known for its efficiency and competitive performance on long-sequence forecasting tasks.
To assess the capabilities of large-scale pretrained models, we consider TimesFM Das et al. (2024), evaluated in a zero-shot inference setting without fine-tuning, thereby highlighting the model’s generalization ability. MOIRAI Woo et al. (2024) is excluded from our evaluation as it is primarily designed for probabilistic forecasting, whereas our setup focuses on point forecasting.
Additionally, we evaluate against Simple-Mamba (S-Mamba) Wang et al. (2025), a recent selective state-space model that leverages a lightweight architecture to efficiently capture both inter-variate correlations and temporal dependencies with near-linear computational complexity.
We also include Repeat Last, a naive baseline that simply repeats the last observed value. This serves as a reference for assessing the forecasting performance of other models.
Classical statistical models such as ARIMA and seasonal ARIMA (SARIMA) are excluded from direct comparisons due to their high computational cost for parameter selection and their consistently poor performance on long sequence time-series forecasting tasks, as reported in prior studies Wu et al. (2021); Zhou et al. (2021). Moreover, we do not consider models that have been specifically designed for a particular domain or dataset (e.g. air pollution or carbon intensity) as part of our baseline comparisons. This is because they cannot be readily used in other domains, often requiring changes to their architecture or pre-processing steps.
5 Results